f(z)= -3z + 5 find the value for f (z + p)

Answer by jim_thompson5910(35256)

Replace every 'z' with 'z+p'


f(z) = -3z + 5
f(z+p) = -3(z+p) + 5
f(z+p) = -3z-3p + 5